Need assistance solving problem.
Question 3 What is the proportion of scores in a normal distribution between z--0.19 and z+3.027 3.02 Report your answer at four decimal places. Moving to another question will save this response.
Answer #1

We have to find P(-0.19<Z<3.02)

P(-0.19<Z<3.02)= P(-0.19<Z<0) + P(0<Z<3.02)

Since Z-Score (Standard Normal distribution is symmetrical about 0)

P(-0.19<Z<3.02)= P(0<Z<0.19) + P(0<Z<3.02) = 0.07534543 + 0.4987361= 0.5740816 ~ 0.57

Here Probability can be calculated by Z-table or in R.

> pnorm(0.19)-0.5
[1] 0.07534543
> pnorm(3.02)-0.5
[1] 0.4987361